OpenStudy (anonymous):

a rectangle parking lot has a lenth that is 7 yards greater than the width. the area of the parking lot is 330 . find the lenth and width.

OpenStudy (anonymous):

let width be x \[length = x+7\] \[x(x+7)=330 \] \[x^2+7x-330=0 \] \[x^2+22x-15x-330 =0 \] \[x(x=22)-15(x=22)=0 \] \[(x-22)(x-15)=0 \] \[x=-22\] or \[x= 15\] ignore negative value of 22 The width is 15 yards. the length will be 15+7 = 22 yards